OpenStudy (anonymous):

4/6 cup of flour .the recipe calls for 1/4 cup of flour .how much flour would you have left if you made the bread would i use subtraction ?

OpenStudy (anonymous):

yes...use subtraction, but make sure you get a common denominator for the two fractions before subtracting.

OpenStudy (anonymous):

also, you might want to reduce 4/6...4/6 = 2/3

OpenStudy (anonymous):

so its now 4/6- 1/4=whats the common demoninator ?

OpenStudy (anonymous):

reducing first, you have 2/3 - 1/4. What is the smallest number that is divisible by both 3 and 4?

OpenStudy (anonymous):

yes, you would use subtraction. you need the get the least common denominator, which in this case is 12. (4/6)2 and (1/4)3. you should end up with (8/12)-(3/12) you'll end up with 5/12

OpenStudy (anonymous):

did that help?

OpenStudy (anonymous):

so 12 is the lcd

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes that did help

OpenStudy (anonymous):

is 2/4+1/5=14/20?

OpenStudy (anonymous):

14/20 can reduce. if you start out with 2/4 = 1/2, you don't need to redduce in the end. 2/4 + 1/5 = 1/2 + 1/5 = 5/10 + 2/10 = 7/10

OpenStudy (anonymous):

Your answer of 14/20 was correct...but not finished. In my opinion, getting the correct answer as you did is the hardest part :)